Suppose a block with a mass of 2.50 kg is resting on a ramp. If the coef-

ficient of static friction between the block and ramp is 0.350, what maximum angle can

the ramp make with the horizontal before the block starts to slip down?


1
Expert's answer
2021-09-24T09:22:57-0400


F ="2.50 \times9.8" = 24.5N


Fp ="24.5\times sinA =" Force parallel with plane.


Fv ="24.5\times cosA" = Force perpendicular to

plane.


uFv"= 0.35F_v =" Force of friction.


Fp - 0.35Fv = ma,

a = 0(block at rest),

Fp -0.35Fv = 0,

Fp = 0.35Fv,

"24.5\times sinA = 0.35\times 24.5\times cosA,"

Divide both sides by 24.5:

sinA = "0.35\times cosA"

Divide both sides by cosA:

"sinA \/ cosA" = 0.35,

"sinA \/ cosA" = TanA,

TanA = 0.35,


maximum angle is

"A = 19.3\degree"